OpenStudy (anonymous):

Find a polynomial function that has the given zeros: (1 + sq. root 3, 1 - sq. root 3)

OpenStudy (anonymous):

2 zeroes means , it is a polynomial of degree 2 , i.e. , it is quadratic polynomial.

OpenStudy (anonymous):

the general form of a quadratic polynomial- \[x^{2}-(\alpha +\beta)x+\alpha \beta=0\] where \[\alpha\]and\[\beta\]are the zeroes/roots of the polynomial. Here, \[\alpha = 1+\sqrt{3}\]\[\beta = 1-\sqrt{3}\] so substitute the values of alpha an beta in the general form to get the desired polynomial.